# Diffusion-Aided Bandwidth-Efficient Semantic Communication with Adaptive Requests

### v3

@@ III-A Simulation Settings and Benchmarks
III-B Experiment Results and Discussions
Figure 2: System performance over different $\tau$ values. (a) Total transmission rounds $t$. (b) Compression ratio $\kappa$. (c) Round-wise histogram.
(a) ROUGE-L.
(b) CLIP-IT.
(c) SSIM.
(d) LPIPS.
(e) FID over SNR.
(f) FID over $\tau$.
Figure 3: System average performance. (a)(b)(c)(d) metrics over SNR values at $\tau=0.7$. (e)(f) FID performance over SNR and $\tau$ values.
Fig. 2 studies the effect of the ROUGE-L threshold $\tau$ on the adaptive loop over all SNR values. As $\tau$ increases, the system becomes more conservative in stopping, which raises the total number of transmissions $t$ by allocating additional retransmission rounds beyond the first pass. Correspondingly, the overall compression ratio $\kappa$ increases with $\tau$ while the drop rate $d$ decreases, reflecting that stricter semantic consistency induces the receiver to request more latent evidence. The round-wise histogram further shows that small $\tau$ concentrates terminations in the first one or two rounds, whereas large $\tau$ shifts mass to later rounds; nevertheless, most sessions still finish within a small $t$ because inpainting effectively propagates sparse evidence across the masked regions. This validates that $\tau$ acts as a direct semantic knob on the rate-fidelity tradeoff, where lower $\tau$ yields aggressive savings through earlier stops and higher $\tau$ spends more bits to drive stronger agreement with the intended description.
Fig. 3(a)-(d) demonstrate system performance under $\tau=0.7$ for two $l$ sizes, comparing the main scheme with the no-guidance scheme. ROUGE-L in Fig. 3(a) remains highest for the main scheme across SNR values, and is relatively insensitive to channel quality. CLIP-IT in Fig. 3(b) also favors the main scheme with small fluctuations over SNR, and prefers $l=8$ over $l=4$ under the main scheme, suggesting that coarser blocks convey more coherent mid-level structure per request that benefits cross-modal matching. On pixel and perceptual metrics, the pattern partly reverses. For SSIM in Fig. 3(c), the main scheme is slightly higher than no-guidance when $l=4$, but slightly lower when $l=8$. For LPIPS in Fig. 3(d), the two methods are nearly identical when $l=4$, while the main scheme is clearly better when $l=8$ because lower LPIPS is preferred. Both SSIM and LPIPS show that $l=4$ performs better than $l=8$, and both slightly worsen as SNR increases. These trends are consistent with a semantic-pixel tradeoff. Caption guidance steers inpainting toward text-aligned content and perceptual coherence, which improves LPIPS values, and it can reduce local pixel agreement in some settings, which affects SSIM and is more pronounced with coarser blocks ($l=8$). Finer blocks enable more targeted fixes that improve local structure and texture statistics, which explains the advantage of $l=4$ for SSIM and LPIPS, even when cross-modal alignment prefers $l=8$ in CLIP-IT. The mild degradation of SSIM and the slight rise in LPIPS with increasing SNR can be explained by the stopping rule. At higher SNR the receiver reaches the ROUGE-L threshold earlier and requests fewer additional blocks, leading to less latent evidence and fewer denoising steps. As a result, pixel-level and perceptual gains are slightly smaller while semantic alignment remains stable.
Fig. 3(e) and (f) analyze distributional realism using FID. In Fig. 3(e), FID becomes slightly worse as SNR increases for all configurations, which is consistent with earlier stopping at high SNR that reduces the amount of corrective evidence the receiver acquires. In Fig. 3(f), increasing the threshold $\tau$ clearly improves FID in our main scheme, since higher $\tau$ triggers additional requests that carry useful information, with diminishing returns at large $\tau$ comparing to the no-guidance scheme. In both subfigures $l=4$ is better than $l=8$, which suggests that finer grained additions align global feature statistics more effectively even when CLIP-IT prefers coarser blocks. Taken together, these results show that the semantic controller yields stable language-level alignment across channels, while the choice of $l$ trades cross-modal alignment (which favors $l=8$) against local and distributional fidelity (which favors $l=4$). Adjusting $\tau$ tunes the balance between realism and rate, with higher $\tau$ improving realism at the expense of more rounds $t$ and lower $\tau$ reducing the number of rounds $t$ and the rate.
Fig. 4 depicts qualitative reconstructions under three compression ratios. Within each row, the caption-only full-mask baseline produces globally plausible but semantically unstable images, with noticeable losses of fine structure at all budgets. The no-guidance scheme tends to preserve local textures more faithfully, yielding higher pixel similarity but exhibiting semantic drift in content and layout, especially at lower budgets. In contrast, the proposed main scheme restores the intended semantics more reliably while maintaining coherent appearance; at $\kappa=0.125$ it already corrects notable mismatches relative to the no-guidance scheme, and as $\kappa$ increases to $0.25$ and $0.375$, a few retransmissions progressively resolve residual inconsistencies without resorting to dense transmission. These visual trends are consistent with the quantitative results, as semantic guidance trades a small amount of pixel-level accuracy for stronger alignment with the intended description and larger $\kappa$ tightens both semantic and perceptual fidelity across methods.
(a) Refer to caption Refer to caption Refer to caption Refer to caption
(b) Refer to caption Refer to caption Refer to caption Refer to caption
(c) Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Reconstructions at termination for sessions with final compression ratio (a) $\kappa=0.125$, (b) $\kappa=0.25$, (c) $\kappa=0.375$, under SNR 77 dB, $l=4$ and $\tau=0.7$. Columns from left to right: original images, full-mask scheme, no-guidance scheme, main scheme.
Finally, we present a brief complexity note. The total compute scales with the sum of denoising steps executed across rounds in the latent diffusion sampler. VAE encode and decode and ROUGE-L scoring add only small overheads relative to denoising. Memory is dominated by U-Net activations and model parameters and remains essentially constant across rounds. In our setup with an RTX 5080 GPU with an AMD 9700X CPU, single-image inference completes in under three seconds.
@@ IV Conclusion
This paper presents a receiver-driven semantic retransmission framework that sends a short caption with a sparse set of latent blocks at the transmitter, and reconstructs images by caption-guided diffusion inpainting at the receiver. A language-based score drives the loop, so the receiver requests additional latent blocks only when the semantic criterion is not met. This strategy aligns the control signal with semantics, reduces the first-pass payload, and spends extra bits where they provide the largest semantic gain. Experiments show substantial bandwidth savings over pixel-based and dense latent baselines while maintaining stable ROUGE-L and consistent perceptual quality. We also observe a trade-off between block size and semantic fidelity, and a balance between image reconstruction and rate control by tuning the threshold. Future work includes learning a block selection policy, modeling reliability for text and latent streams, and extending the framework to video and multi-user scenarios.

### v4

@@ III-A Simulation Settings and Benchmarks
III-B Experiment Results
Figure 2: Effect of the ROUGE-L threshold $\tau$. (a) Total number of transmission rounds (including the first pass and feedback rounds). (b) Final block sending rate $\kappa$. (c) Histogram of termination round index $t$.
(a) ROUGE-L
(b) CLIP-IT
(c) SSIM
(d) LPIPS
(e) FID over SNR
(f) FID over $\tau$
Figure 3: System performance. (a)(b)(c)(d) metrics over SNR values at $\tau=0.7$. (e)(f) FID performance over SNR and $\tau$ values.
Fig. 2 shows that the threshold $\tau$ directly controls how conservative the closed loop is in early stopping. As $\tau$ increases, the receiver requires stronger semantic agreement before stopping, so the termination round index shifts to later rounds and the average number of rounds increases. This increases the total number of requested latent blocks and raises the sending ratio $\kappa$, i.e., more semantic content is ultimately delivered. The histogram further indicates that most sessions terminate within a few rounds for all tested $\tau$, while larger $\tau$ mainly increases the probability of late termination.
Fig. 3(a)-(d) report the average performance at $\tau=0.7$ for two block sizes $l\in\{4,8\}$, comparing the main scheme with four baselines. On semantic-oriented metrics, the main scheme achieves the highest ROUGE-L and is among the top CLIP-IT performers across SNR values, with fixed-budget being close on CLIP-IT. The no-guidance baseline (latent-only) degrades substantially, indicating that without a textual constraint the generative prior can drift in high-level content even when partial latent evidence is available. The full-mask (caption-only) baseline lacks source-specific latent evidence, leading to the weakest pixel/perceptual quality and poorer semantic alignment among the baselines. Notably, while fixed-round scheme runs more rounds, it does not necessarily increase the caption-consistency score. Each round re-runs stochastic diffusion inpainting with additional revealed blocks, and the sampler is not designed to monotonically improve a language-based criterion, so the caption-consistency score is not guaranteed to improve monotonically, since diffusion-based inpainting is stochastic and can yield different plausible completions even when more evidence is revealed [10].
Comparing the main scheme with the fixed-budget one-shot baseline isolates the benefit of receiver-driven adaptation under the same average latent payload. Although the fixed-budget baseline transmits $N_{\mathrm{fix}}$ blocks matched to the main scheme on average, it cannot adapt its payload to sample difficulty and therefore yields lower ROUGE-L. This is the key advantage of the closed loop. It spends extra blocks only when the receiver-side semantic criterion is not met, while stopping early on easier samples.
On pixel/perceptual fidelity (SSIM/LPIPS), the fixed-round scheme performs best because it always runs $t_{\max}$ rounds and thus accumulates substantially more retransmitted latent blocks than the early-stopping schemes (as shown in Fig. 2(a), where the main scheme typically terminates after far fewer rounds). This curve should therefore be interpreted as a high-overhead reference rather than a rate-comparable baseline. In contrast, the main and fixed-budget schemes are close on SSIM/LPIPS, indicating that matching the average latent payload largely closes the pixel/perceptual gap; the main scheme primarily differs by reallocating the same budget across images via semantic-driven stopping.
Fig. 3(e) shows that FID mildly increases with SNR for some configurations. This is consistent with our heuristic that decreases the inpainting strength and therefore the number of denoising steps at higher SNR. The sampler relies more on the received latent evidence and performs fewer corrective diffusion updates, which can slightly worsen dataset-level feature statistics. Fig. 3(f) further shows that increasing $\tau$ improves FID for the main and no-guidance schemes because more evidence is delivered before stopping, while the caption-only full-mask baseline degrades as $\tau$ increases.
Fig. 4 reports $P_{\text{fail}}$ versus SNR at $\tau=0.7$. The main scheme consistently achieves the lowest failure probability, since feedback injects additional latent evidence only when the semantic score is insufficient. Fixed-round can show higher $P_{\text{fail}}$ despite sending more blocks, because the caption-consistency score is not guaranteed to improve monotonically under stochastic diffusion re-inpainting [10].
Fig. 5 shows a representative example at $\kappa=0.25$. The full-mask baseline produces a plausible image but lacks source-specific details, while the no-guidance baseline tends to preserve local textures yet may drift in high-level semantics. In contrast, the main scheme progressively refines the reconstruction across rounds as additional latent blocks are requested, and the reconstruction becomes increasingly consistent across rounds, reducing inconsistent regions while keeping the content aligned with the caption constraint.
Figure 4: Semantic failure probability $P_{\text{fail}}$ versus SNR at $\tau=0.7$.
(a) Refer to caption (b) Refer to caption (c) Refer to caption
(d) Refer to caption (e) Refer to caption (f) Refer to caption
Figure 5: Qualitative example at $\kappa=0.25$ (SNR =7=7 dB, $l=4$, $\tau=0.7$): (a) original, (b) full-mask, (c) no-guidance at $t=2$, (d)-(f) main at $t=0,1,2$.
Finally, we comment on complexity. Runtime is dominated by diffusion sampling, and the overall cost scales with the cumulative number of denoising steps across rounds, i.e., $\sum_{t}S_{t}$. VAE encode/decode, caption generation, and score computation run once per round and add a small constant overhead relative to sampling. Memory is dominated by the fixed model parameters and U-Net activations during sampling, and thus remains essentially constant across retransmission rounds.
@@ IV Conclusion
This paper proposes a receiver-driven semantic retransmission framework for image transmission with a generative receiver. The transmitter sends a short caption and an initial sparse subset of VAE latent blocks, while the receiver performs caption-conditioned latent diffusion inpainting and iteratively requests additional blocks until a receiver-side semantic criterion is met or a round limit is reached. Experiments on Flickr30k over AWGN channels confirm a tunable rate-quality tradeoff. Compared with budget-matched one-shot sparse transmission, closing the loop yields higher ROUGE-L semantic consistency and fewer semantic failures, and it generally requires fewer latent blocks than an always-on retransmission scheme. Overall, the results indicate that receiver-side semantic control can improve reliability for generative receivers under a fixed average evidence budget. Future work will consider importance-aware block selection, unreliable captions/feedback, and extensions to video and multi-user settings.